# Hi-Core: Hierarchical Knowledge Transfer for Continual Reinforcement Learning

### v1

@@ 4 Experiments
In this section, we evaluate our framework in several continual reinforcement learning tasks.
@@ 4.1 Environment
For our experiments, we utilized a suite of MiniGrid environments Chevalier-Boisvert et al. (2023) to evaluate the efficacy of Hi-Core in addressing CRL tasks. These environments feature image-based state observations, a discrete set of possible actions, and various objects characterized by their color and type. We intentionally simplified the state and action spaces to streamline the policy learning process, enabling more rapid experimental iterations. This approach allowed us to focus on the core objective of this study: leveraging hierarchical knowledge transfer for CRL. We crafted a sequence of four distinct tasks within the MiniGrid framework. Visual representations of each task environment are provided in Figure 3. This sequence is arranged in ascending order of difficulty and aligns with the human learning process, facilitating better tracking of knowledge transfer throughout the learning process.
@@ 4.2 Experiment Setup
Implementation. We use OpenAI’s GPT as the LLM, which can be accessed through their API. To improve the language model’s reasoning ability, we introduce the Chain of Thought (COT) Wei et al. (2022) in the prompt. This approach also aids the LLM in addressing complex tasks. Moreover, we add a few-shot example at the end of the prompt to guide the output towards the desired format. Note that this example contains only the foundational knowledge of all tasks, requiring the LLM to engage in additional reasoning to comprehend the specifics of the current task. For the low-level policy learning in Hi-Core, we chose the Proximal Policy Optimization (PPO) algorithm as the underlying RL algorithm due to its excellent and stable performance Schulman et al. (2017); Raffin et al. (2021). For the policy library, we build a vector database based on FAISS, which is a library for efficient similarity search and clustering of dense vectors Johnson et al. (2019). Due to the small total number of tasks, the number of experiences to be retrieved can be set as the total number of tasks . In addition, the policy library in our experiments does not store low-level policies for fair comparison with limited storage space.
Baselines. We consider the following baselines, for comparison with our approach in experiments:
Single-Task (SG): This baseline represents a traditional RL setup where a distinct agent is trained exclusively on each task. This serves as a foundational comparison point to underscore the advantages of CL, as it lacks any mechanism for knowledge retention or transfer.
Fine-Tuning (FT): Building upon the standard RL algorithm, this baseline differs from SG by using a single agent that is sequentially fine-tuned across different tasks. This baseline provides a basic measure of an agent’s capacity to maintain knowledge of earlier tasks while encountering new tasks.
Note that our method is not directly comparable with certain methods, such as CLEAR Rolnick et al. (2019) and ClonEx-SAC Wolczyk et al. (2022), as they require access to the data from previous tasks for training. Our framework can continually learn without the need to revisit past task data, which is a common constraint in real-world applications.
Metrics. To evaluate the effectiveness of Hi-Core, we use the normalized average return to measure the performance of the trained agents. Following standard practice in supervised continual learning literature Díaz-Rodríguez et al. (2018); Wolczyk et al. (2021), we define a suite of metrics based on the agent’s performance throughout different phases of its training process. These metrics include average performance , forward transfer and forgetting .
@@ 4.3 Competitive Results
To evaluate our framework, we compare Hi-Core with other baselines across four challenging tasks that have the same state and action space (Figure 3). The underlying RL algorithm of all baselines is PPO. Each experiment is trained in 5M steps and replicated with five random seeds of environments to ensure reliability.
Overall Performance. As illustrated in Figure 4, the baselines perform well on the first Task. However, their performance gets worse progressively as the difficulty of the tasks increases. A common trend of performance degradation in the transition between tasks is evident across all methods. Specifically, the SG baseline, while effective on the first task, is unable to maintain its performance due to its inherent limitation. Other baselines, such as FT, FT-L2, and PackNet, also start strong on the first task but encounter difficulties in the subsequent tasks. This phenomenon suggests that a singular shared policy may not be robust enough to handle a sequence of highly varied tasks. The balance between policy stability and plasticity is critical, with higher stability often leading to reduced adaptability, as observed in the performance degradation of these baselines. In contrast, Hi-Core consistently surpasses the other baselines, showcasing remarkable performance across all tasks. Although it is not immune to performance degradation when transitioning between tasks with significant differences, it demonstrates a rapid recovery in performance. This can be attributed to its hierarchical structure, which leverages the powerful capabilities of LLM and promotes the transfer of high-level policy knowledge, thereby enhancing the agent’s ability to adapt to new and diverse tasks.
| Metric | SG | FT | FT-L2 | PackNet | Hi-Core |
@@
Continual Learning Performance. Table 1 shows the evaluation results in terms of CL metrics. The average performance metric clearly demonstrates Hi-Core’s superiority in task sequence with diverse tasks, significantly outperforming other baseline methods. The forward transfer metric is particularly important, as it measures the capacity of an agent to utilize knowledge from previous tasks. All baselines, with the exception of Hi-Core, present forward transfer metrics of less than or equal to zero. This suggests that traditional approaches struggle to transfer previously acquired knowledge to new and diverse tasks effectively. Baselines that rely on regularization strategies, such as FT-L2, or parameter isolation techniques, like PackNet, may even hinder the learning of new tasks.
In contrast, Hi-Core exhibits robust transfer performance, underscoring the benefits of transferring high-level policies. When examining the forgetting metrics, FT, FT-L2, and PackNet show relatively low scores. This can be attributed, in part, to their emphasis on model stability. However, a more critical factor is their inherently lower performance on subsequent tasks, such as Task 2, Task 3, and Task 4. Although Hi-Core does not achieve the lower score in the forgetting metric, its exceptional forward transfer capability and strong average performance accentuate its proficiency in handling diverse task sequences. Note that Hi-Core does not resort to fine-tuning strategy in this experiment, although it could be employed in situations where minimizing forgetting is a priority. The above findings highlight Hi-Core’s potential to markedly enhance the adaptability and longevity of reinforcement learning agents, positioning it as a highly viable solution for dynamic and diverse tasks.
@@
| Hi-Core | 0.65 | 0.17 | |
| Hi-Core-FT | 0.09 | ||
| Hi-Core-SG | |||
| Hi-Core-Once |
4.4 Heterogeneous Tasks Experiments
Most existing works have focused on the transfer ability of CRL agents in environments that have the same state space. In contrast, our study explores the application of CRL agents to heterogeneous tasks, where environments have different state spaces or action spaces. For the simplification of the problem, we only concentrate on tasks with differing state spaces.
As depicted in Figure 5a, heterogeneous tasks present a unique challenge. While it is possible to accommodate differences in the size of the state space during knowledge transfer by employing a masking technique, this approach necessitates prior knowledge of the maximum state space size, which is impractical for RL agents expected to learn continually in dynamic environments. On the contrary, the high-level policy transfer of Hi-Core can be applied to heterogeneous tasks without the need for masking.
In our experiments, we evaluate the effectiveness of Hi-Core across two heterogeneous tasks. The first task has a relatively small state space, as shown in the left below of Figure 5a. In contrast, the second task has a larger state space, such as the right below of the same figure. Figure 5b reports the results of these experiments. The first task is Task 1, while the second task is Task L1, Task L2, Task L3, or Task L4, each representing a larger state space variant of Task 1 through Task 4. Direct learning agents who are exposed to tasks with large state spaces from the outset may struggle to learn effective policies within the set number of steps, resulting in poor performance. This phenomenon is more obvious on more difficult tasks, such as Task L3 and Task L4. However, agents that are learned on Task 1 can transfer high-level knowledge to subsequent tasks. Despite the increasing difficulty of these tasks, the agents still demonstrate comparable performance. This evidence suggests that Hi-Core retains transfer ability even when applied to heterogeneous tasks.
4.5 Ablation Study
@@
In this part, we conduct an ablation study to investigate what affects Hi-Core’s performance in CRL. We consider several variants of Hi-Core to understand the effects:
@@
Hi-Core-FT: Hi-Core fine-tuning the policy from the previous task when encountering a new task.
@@
Hi-Core-SG: Hi-Core without policy library. This variant treats each task as an isolated learning problem, akin to training single agents for each task.
@@
Hi-Core-Once: Hi-Core without feedback mechanisms. In this variant, the high-level policy formulation of LLM is performed only once for each task.
@@
The results of our study are presented in Table 2 and Figure 6. For better comparison, we also include the performance of the SG baseline in the Table. Our findings are as follows: Firstly, Hi-Core-SG shows superior performance compared to the SG baseline, which confirms that the LLM’s powerful capabilities can help RL. An RL agent can take advantage of reasoning ability in the LLM to significantly reduce the difficulties in solving these tasks. Secondly, while Hi-Core-SG outperforms the baseline, it falls short of the average performance and forward transfer metrics achieved by the Hi-Core. As shown in Figure 6b, this gap emphasizes the value of the policy library in facilitating knowledge transfer, ultimately boosting the agent’s overall performance. Furthermore, Hi-Core-Once exhibits slightly lower performance on Task 2, Task 3 and Task 4 than Hi-Core. This phenomenon suggests that feedback can improve the performance of LLM in dealing with these challenging tasks. Lastly, the Hi-Core-FT variant underperforms on all tasks except Task 1. This observation aligns with the previous finding in competitive results that methods designed to enhance policy stability may decrease the agent’s performance on subsequent tasks when those tasks differ significantly from one another.
@@
Policy Visualization. Figure 7 further illustrates the difference between Hi-Core and Hi-Core-SG on Task 2 and Task 4. As shown in Figure 7a, the agent of Hi-Core demonstrates the capability to swiftly learn to pick up the blue key and unlock the corresponding door. In contrast, the agent of Hi-Core-SG struggles to progress beyond the initial room (Figure 7b). This performance gap can be attributed to the previous learning of the Hi-Core agent on Task 1, which shares common elements with Task 2, such as the blue key and door. By drawing on previous experience, Hi-Core formulate relevant initial goals, accelerating learning on Task 2. Similarly, when comparing Figure 7c and Figure 7d for Task 4, we observe a consistent pattern. Although Task 4 presents a higher level of difficulty, the agent with pre-acquired high-level knowledge exhibits more effective behaviors than the agent learning from scratch. These findings demonstrate the effectiveness of Hi-Core in facilitating the transfer of knowledge across diverse tasks.
@@ 5 Conclusion
In this work, we have presented Hi-Core, a novel framework that aligns with the hierarchical structure of human cognitive control, enhancing the high-level knowledge transfer of CRL. Hi-Core leverages a LLM for high-level policy formulation, employs goal-oriented RL for low-level policy learning, and constructs the policy library to store and retrieve policies. It addresses the challenge of learning across a sequence of diverse tasks by transferring high-level knowledge, which has been demonstrated in our experiments. Further results indicate that Hi-Core can be applied to heterogeneous tasks and achieves comparable transfer performance.
We believe that this work constitutes the first step toward the use of hierarchical knowledge transfer to enhance the effectiveness and generalization of CRL agents. There are still some limitations of Hi-Core, such as an unsatisfactory solution to catastrophic forgetting and room for improvement in the transferability of low-level policies. These directions provide opportunities for future work, with the potential to refine and extend Hi-Core.

### v3

@@ 4 Experiments
In this section, we evaluate our framework in several continual reinforcement learning tasks. More details about the experimental implementation are provided in the supplementary material.
@@ 4.1 Environment
For our experiments, we utilized a suite of MiniGrid environments Chevalier-Boisvert et al. (2023) to evaluate the efficacy of MT-Core in addressing CRL tasks. These environments feature image-based state observations, a discrete set of possible actions, and various objects characterized by their color and type. We intentionally simplified the state and action spaces to streamline the policy learning process, enabling more rapid experimental iterations. This approach allowed us to focus on the core objective of this study: leveraging hierarchical knowledge transfer for CRL. We crafted a sequence of four distinct tasks within the MiniGrid framework. This sequence is arranged in ascending order of difficulty and aligns with the human learning process, facilitating better tracking of knowledge transfer throughout the learning process. A comprehensive account of these modified environments, including their specific configurations and visual representations, is available in supplementary material.
@@ 4.2 Experiment Setup
Baselines.
We consider the following baselines:
Single-Task (SG): This baseline represents a traditional RL setup where a distinct agent is trained exclusively on each task. There is no sharing or transfer of knowledge between tasks in this method. This serves as a foundational comparison point to underscore the advantages of CL, as it lacks any mechanism for knowledge retention or transfer.
Fine-Tuning (FT): Building upon the standard RL algorithm, this baseline differs from SG by using a single agent that is sequentially fine-tuned across different tasks. The agent trained on the preceding task is adapted for the subsequent task, effectively using the prior model as a starting point. This baseline provides a basic measure of an agent’s capacity to maintain knowledge of earlier tasks while encountering new tasks.
Fine-Tuning with L2 Regularization (FT-L2): Based on FT, this baseline incorporates L2 regularization during the fine-tuning process. The addition of L2 regularization aims to mitigate catastrophic forgetting by penalizing significant changes to the weights that are important for previous tasks. This helps in preserving performance on earlier tasks while learning new ones. FT-L2 can be viewed as the simplest implementation of regularization-based CRL methods Kirkpatrick et al. (2017); Wang et al. (2024c).
PackNet Mallya and Lazebnik (2018): This baseline is a representative CRL method based on parameter isolation Wang et al. (2024c). It utilizes the network pruning technique to efficiently allocate the neural network’s capacity throughout the continual learning of tasks. After training on a task, PackNet identifies the crucial weights to retain, prunes the less essential ones, and thus creates room for the network to learn additional tasks. PackNet is a representative CRL method based on parameter isolation Wang et al. (2024c).
Note that our method is not directly comparable with certain methods, such as CLEAR Rolnick et al. (2019) and ClonEx-SAC Wolczyk et al. (2022), as it requires access to the data from previous tasks for training. Our framework can continually learn without the need to revisit past task data, which is a common constraint in real-world applications.
Metrics.
To evaluate the effectiveness of MT-Core, we use the normalized average return to measure the performance of the trained agents. Following standard practice in supervised continual learning literature Wolczyk et al. (2021); Li et al. (2024), we define a suite of metrics based on the agent’s performance throughout different phases of its training process. Based on the agent’s normalized average return, we evaluated the continual learning performance of our framework and baselines using the following metrics: average performance, forward transfer and forgetting, as detailed in the supplemental material.
@@ 4.3 Competitive Results
To evaluate our framework, we compare MT-Core with other baselines across four challenging tasks that have the same state and action space. Each experiment is trained in 5M steps and replicated with five random seeds of environments to ensure reliability.
Overall Performance. As illustrated in Figure 3, the baselines perform well on the first Task. However, their performance gets worse progressively as the difficulty of the tasks increases. A common trend of performance degradation in the transition between tasks is evident across all methods. Specifically, the SG baseline, while effective on the first task, is unable to maintain its performance due to its inherent limitation. Other baselines, such as FT, FT-L2, and PackNet, also start strong on the first task but encounter difficulties in the subsequent tasks. This phenomenon suggests that a singular shared policy may not be robust enough to handle a sequence of highly varied tasks. The balance between policy stability and plasticity is critical, with higher stability often leading to reduced adaptability, as observed in the performance degradation of these baselines. In contrast, MT-Core consistently surpasses the other baselines, showcasing remarkable performance across all tasks. Although it is not immune to performance degradation when transitioning between tasks with significant differences, it demonstrates a rapid recovery in performance. This can be attributed to its multi-granularity structure, which leverages the powerful capabilities of LLM and promotes the transfer of coarse-grained policy knowledge, thereby enhancing the agent’s ability to adapt to new and diverse tasks.
| Metric | SG | FT | FT-L2 | PackNet | MT-Core |
@@
Continual Learning Performance. Table 1 shows the evaluation results in terms of CL metrics. The average performance metric clearly demonstrates MT-Core’s superiority in task sequence with diverse tasks, significantly outperforming other baseline methods. The forward transfer metric is particularly important, as it measures the capacity of an agent to utilize knowledge from previous tasks. All baselines, with the exception of MT-Core, present forward transfer metrics of less than or equal to zero. In addition, the average performace of MT-Core is more than twice that of the best baseline. This suggests that traditional approaches struggle to transfer previously acquired knowledge to new tasks effectively. Baselines that rely on regularization strategies, such as FT-L2, or parameter isolation techniques, like PackNet, may even hinder the learning of new tasks.
In contrast, MT-Core exhibits positive transfer performance, underscoring the benefits of transferring coarse-grained policies. When examining the forgetting metrics, FT, FT-L2, and PackNet show relatively low scores. This can be attributed, in part, to their emphasis on model stability. However, a more critical factor is their inherently lower performance on subsequent tasks, such as Task 2, Task 3, and Task 4. Although MT-Core does not achieve the lower score in the forgetting metric, its exceptional forward transfer capability and strong average performance accentuate its proficiency in handling diverse task sequences. Note that MT-Core does not resort to fine-tuning strategy in experiments to focus on coarse-grained knowledge transfer, although it could be employed in situations where minimizing forgetting is a priority.
4.4 Heterogeneous Tasks Experiments
Most existing works have focused on the transfer ability of CRL agents in environments that have the same state space. In contrast, our study explores the application of CRL agents to heterogeneous tasks, where environments have different state spaces or action spaces. For the simplification of the problem, we concentrate on tasks with differing state spaces.
As depicted in Figure 4a, heterogeneous tasks present a unique challenge. While it is possible to accommodate differences in the size of the state space during knowledge transfer by employing a masking technique, this approach necessitates prior knowledge of the maximum state space size, which is impractical for RL agents expected to learn continually in dynamic environments. On the contrary, the coarse-grained policy transfer of MT-Core can be applied to heterogeneous tasks without the need for masking.
In our experiments, we evaluate the effectiveness of MT-Core across two heterogeneous tasks. The first task has a relatively small state space, as shown in the left below of Figure 4a. In contrast, the second task has a larger state space, such as the right below of the same figure. Figure 4b reports the results of these experiments. The first task is Task 1, while the second task is Task L1, Task L2, Task L3, or Task L4, each representing a larger state space variant of Task 1 through Task 4. Direct learning agents who are exposed to tasks with large state spaces from the outset may struggle to learn effective policies within the set number of steps, resulting in poor performance. This phenomenon is more obvious on more difficult tasks, such as Task L3 and Task L4. However, agents that are learned on Task 1 can transfer coarse-grained knowledge to subsequent tasks. Despite the increasing difficulty of these tasks, the agents still demonstrate comparable performance. This evidence suggests that MT-Core retains transfer ability even when applied to heterogeneous tasks.
4.5 Ablation Study
@@
| MT-Core | 0.65 | 0.17 | |
| MT-Core-FT | 0.09 | ||
| MT-Core-SG | |||
| MT-Core-Once |
@@
In this part, we conduct an ablation study to investigate what affects MT-Core’s performance in CRL. We consider several variants of MT-Core to understand the effects:
@@
MT-Core-FT: MT-Core fine-tuning the policy from the previous task when encountering a new task.
@@
MT-Core-SG: MT-Core without policy library. This variant treats each task as an isolated learning problem, akin to training single agents for each task.
@@
MT-Core-Once: MT-Core without feedback mechanisms. In this variant, the coarse-grained policy formulation of LLM is performed only once for each task.
@@
The results of our study are presented in Table 2 and Figure 5. For better comparison, we also include the performance of the SG baseline in the Table. Our findings are as follows: Firstly, MT-Core-SG shows superior performance compared to the SG baseline, which confirms that the LLM’s powerful capabilities can help RL. An RL agent can take advantage of reasoning ability in the LLM to significantly reduce the difficulties in solving these tasks. Secondly, while MT-Core-SG outperforms the baseline, it falls short of the average performance and forward transfer metrics achieved by the MT-Core. The gap between MT-Core with MT-Core-SG emphasizes the value of the policy library in facilitating knowledge transfer, ultimately boosting the agent’s overall performance (see the figure in 5b). Furthermore, MT-Core-Once exhibits slightly lower performance on Task 2, Task 3 and Task 4 than MT-Core. This phenomenon suggests that feedback can improve the performance of LLM in dealing with these challenging tasks. Lastly, the MT-Core-FT variant underperforms on all tasks except Task 1. This observation aligns with the previous finding in competitive results that methods designed to enhance policy stability may decrease the agent’s performance on subsequent tasks when those tasks differ significantly from one another.
@@
Policy Visualization. Figure 6 further illustrates the difference between MT-Core and MT-Core-SG on Task 2 and Task 4. As shown in Figure 6a, the agent of MT-Core demonstrates the capability to swiftly learn to pick up the blue key and unlock the corresponding door. In contrast, the agent of MT-Core-SG struggles to progress beyond the initial room (Figure 6b). This performance gap can be attributed to the previous learning of the MT-Core agent on Task 1, which shares common elements with Task 2, such as the blue key and door. By drawing on previous experience, MT-Core formulate relevant initial goals, accelerating learning on Task 2. Similarly, when comparing Figure 6c and Figure 6d for Task 4, we observe a consistent pattern. Although Task 4 presents a higher level of difficulty, the agent with pre-acquired coarse-grained knowledge exhibits more effective behaviors than the agent learning from scratch.
@@ 5 Conclusion
In this work, we took a step in the direction of improving the generalization ability of CRL agents. We investigated the limitations of existing CRL methods in transferring fine-grained knowledge across diverse tasks and proposed MT-Core, a novel framework that aligns with the multi-granularity structure of human cognitive control, enhancing the coarse-grained knowledge transfer of CRL. MT-Core leverages a LLM for coarse-grained policy formulation, employs goal-oriented RL for fine-grained policy learning, and constructs the policy library to store and retrieve policies. It addresses the challenge of learning across a sequence of diverse tasks by transferring coarse-grained knowledge, which has been demonstrated in our experiments. Further results indicate that MT-Core can be applied to heterogeneous tasks and achieves comparable transfer performance.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Nos. 62476228), the Sichuan Science and Technology Program (Nos. 2024ZYD0180), and the Graduate Representative Achievement Cultivation Project of Southwest University of Finance and Economics (Nos. JGS2024068).